OpenStudy (anonymous):

What is equal to sqrt50a^6 b^7

OpenStudy (ranga):

Separate the expression into perfect squares and non-perfect squares. For example, write 50 as 25*2 because 25 is a perfect square and we take the square root we will get 5. \[\sqrt{50a^6b^7} = \sqrt{2*25*a^6*b^6*b} = \sqrt{2b}*\sqrt{25a^6b^6} = \sqrt{2b}*5a^3b^3\]